OpenStudy (anonymous):

Inverse Functions! Use the theory of inverse functions to prove that f and g are inverse functions of eachother, and sketch the graph of f and g on the same coordinate plane. f(x)=-x^2 + 3, x>=0 g(0)=sqrt(3-x), x<=3 I have no idea how I would start this D: Any help would be amazing! Thank you!!

OpenStudy (ash2326):

Let's first check if they are inverse functions. If g(x) is the inverse of f(x) then \[f(g(x)=g(f(x)=x\] Do you get this?

OpenStudy (ash2326):

@Kilochan ???

OpenStudy (anonymous):

Yes I did get that. But idk what to do after, so make a graph of two functions :O

OpenStudy (anonymous):

Graph the equations as you would normally: g(x) and f(x) within the specified domain for x.

OpenStudy (anonymous):

by "specidied domain" you mean x>=0 and x<=3? I've tried some graphing but they never came out as they were supposed to.

OpenStudy (anonymous):

I will help.

OpenStudy (anonymous):

So visually, from 0 to infinity, f(x) is graphed, and from -infinity to 3, g(x) is graphed. http://www.wolframalpha.com/input/?i=f%28x%29%3D-x%5E2+%2B+3%2C+g%28x%29%3Dsqrt%283-x%29

OpenStudy (anonymous):

Hmmm. I did this exactly and the graph is coming out inverted. idk if I have a wrong mode on my calculator. it's in degree mode. My graphs are upside down. I put "-sqrt(3-x) and "-3-x^2" idk why they arent coming out the same. But your graphs are correct!

OpenStudy (anonymous):

When I get rid of the negative signs, they come out right :O

OpenStudy (agent0smith):

@Kilochan i think you entered the wrong functions. You posted in the original question "f(x)=-x^2 + 3, x>=0 g(0)=sqrt(3-x)," then posted a moment ago "-sqrt(3-x) and "-3-x^2" Google can plot graphs too: https://www.google.com/search?q=sqrt(3-x)%2C+-x%5E2%2B3&oq=sqrt(3-x)%2C+-x%5E2%2B3&aqs=chrome.0.57j62l3.341&sourceid=chrome&ie=UTF-8

OpenStudy (anonymous):

@agent0smith Yes, you are very right. I must have done some type of miscalculation. But now I see where I went wrong! Thank you everyone for your help!